OpenStudy (anonymous):

A company manufacturers and sells x electric drills per month. The monthly cost and price-demand equations are C(x)=69000+40x, p=190−x30, 0≤x≤5000. (A) Find the production level that results in the maximum revenue. Production Level = (B) Find the price that the company should charge for each drill in order to maximize profit. Price = (C) Suppose that a 5 dollar per drill tax is imposed. Determine the number of drills that should be produced and sold in order to maximize profit under these new circumstances. Number of drills =

OpenStudy (anonymous):

A) Revenue: R(x) = P(x) * x R'(x) = 0 => x = .... B) Profit: P(x) = R(x) - C(x) P'(x) = 0 => x = ...